### v1

@@ 5.1 Semi-discrete GePUP-ES-RK schemes
$\mathbf{w}^{(i)}$ $=\mathbf{w}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\rho}^{(j)}$ in $\text{in }\Omega,$ (56d)
$\mathbf{w}^{(i)}\cdot\bm{\tau}$ $=0,~{}\nabla\cdot\mathbf{w}^{(i)}=0$ on $\text{on }\partial\Omega,$ (56e)
@@
$\mathbf{u}^{(i)}$ $=\mathscr{P}\mathbf{w}^{(i)}$ in $\text{in }\Omega,$ (56g)
$\mathbf{n}\cdot\mathbf{u}^{(i)}$ $=0$ on $\text{on }\partial\Omega.$ (56h)
In Definition 6, the approximations $\widetilde{\mathbf{u}}^{(i)}$ in $\Omega$ may be obtained by polynomial interpolation based on stage values of recent time steps. In this work, we fit a cubic polynomial $\mathbf{p}$ from the known velocity $\hat{\mathbf{u}}_{n}^{(j)}$ at time instances $t^{n}+\hat{c}_{j}k$ with $j=0,1,2,3$ and then approximate $\mathbf{u}(t^{n}+c_{i}k)$ with
@@
  1. [(EAU-1)]
  2. 1.
    for the first two time steps $n=0,1$, we calculate $\hat{\mathbf{u}}^{(j)}$ by GePUP-ERK [5], the explicit RK method for solving the GePUP formulation, with time step size $\frac{1}{3}k$ and then fit $\mathbf{p}$ in (57) with $\left[\hat{c}_{0},\hat{c}_{1},\hat{c}_{2},\hat{c}_{3}\right]=\left[0,\frac{1}{3},\frac{2}{3},1\right]$;
  3. 2.
@@
$t_{n}$$t_{n+1}$$\widetilde{\mathbf{u}}_{n}^{\left(1\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(2\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(3\right)}$$t_{n}+c_{1}k$$t_{n}+c_{2}k$$t_{n}+c_{3}k$$t_{n}$$t_{n}+\frac{1}{3}k$$t_{n}+\frac{3}{3}k=t_{n+1}$$\mathbf{u}^{n}$$\mathbf{u}^{n+1}$$\hat{\mathbf{u}}_{n}^{\left(0\right)}=\mathbf{u}^{n}$$\hat{\mathbf{u}}_{n}^{\left(1\right)}$$\hat{\mathbf{u}}_{n}^{\left(3\right)}$$t_{n}+\frac{2}{3}k$$\hat{\mathbf{u}}_{n}^{\left(2\right)}$
(a) (EAU-1) for $n=0,1$
$t_{n}$$t_{n+1}$$t_{n-2}$$\widetilde{\mathbf{u}}_{n}^{\left(1\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(2\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(3\right)}$$t_{n}+c_{1}k$$t_{n}+c_{2}k$$t_{n}+c_{3}k$$t_{n}$$t_{n-1}$$t_{n+1}=t_{n}+k$$\mathbf{u}^{n}$$\mathbf{u}^{n+1}$$\hat{\mathbf{u}}_{n}^{\left(2\right)}=\mathbf{u}^{n}$$\hat{\mathbf{u}}_{n}^{\left(1\right)}=\mathbf{u}^{n-1}$ $\hat{\mathbf{u}}_{n}^{\left(3\right)}$$\hat{\mathbf{u}}_{n}^{\left(0\right)}=\mathbf{u}^{n-2}$
(b) (EAU-2) for $n\geq 2$
Figure 1: Estimating $\widetilde{\mathbf{u}}^{(i)}$ in GePUP-ES-RK by (57) and (EAU-1,2).
Theorem 9.
Suppose the RK method employed in the GePUP-ES-RK scheme (56)–(55) is algebraically stable in the sense of Definition 2. Then (41) implies
@@
$\forall n\in\mathbb{N},\ \forall\mathbf{x}\in\partial\Omega,\quad\left|\mathbf{n}\cdot\mathbf{w}^{n+1}(\mathbf{x})\right|^{2}-\left|\mathbf{n}\cdot\mathbf{w}^{n}(\mathbf{x})\right|^{2}\leq-2k\lambda\sum_{i=1}^{s}b_{i}\left|\mathbf{n}\cdot\mathbf{w}^{(i)}(\mathbf{x})\right|^{2}.$ (59)
Proof.
[…]
@@
$\begin{array}[]{rl}&\left\langle\nabla\cdot\mathbf{w}^{n+1},\nabla\cdot\mathbf{w}^{n+1}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{w}^{n}+k\sum_{i=1}^{s}b_{i}\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\mathbf{w}^{n}+k\sum_{j=1}^{s}b_{j}\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{w}^{n},\nabla\cdot\mathbf{w}^{n}\right\rangle+k\sum_{i=1}^{s}b_{i}\left\langle\nabla\cdot\mathbf{w}^{n},\nabla\cdot\bm{\rho}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\nabla\cdot\mathbf{w}^{n},\nabla\cdot\bm{\rho}^{(j)}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{w}^{n},\nabla\cdot\mathbf{w}^{n}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ &+k\sum_{i=1}^{s}b_{i}\left\langle\nabla\cdot\mathbf{w}^{(i)}-k\sum_{j=1}^{s}a_{i,j}\nabla\cdot\bm{\rho}^{(j)},\nabla\cdot\bm{\rho}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\nabla\cdot\mathbf{w}^{(j)}-k\sum_{i=1}^{s}a_{j,i}\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{w}^{n},\nabla\cdot\mathbf{w}^{n}\right\rangle-2\nu k\sum_{i=1}^{s}b_{i}\left\|\nabla\left(\nabla\cdot\mathbf{w}^{(i)}\right)\right\|^{2}\\ &-k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle,\end{array}$
@@
where the last step follows from (38), Green’s formula (35), and (56e). By Definition 2, the algebraic stability implies the symmetric positive semi-definiteness of $M$ and thus we have $M=O\Lambda O^{T}$, where $O$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix with $\lambda_{\ell}=\Lambda_{\ell,\ell}\geq 0$ for each $\ell=1,\ldots,s$. Hence (58) follows from
@@
$\begin{array}[]{rl}&\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{i=1}^{s}\sum_{j=1}^{s}\left(\sum_{\ell=1}^{s}o_{i,\ell}\lambda_{\ell}o_{j,\ell}\right)\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{\ell=1}^{s}\lambda_{\ell}\left\langle\sum_{i=1}^{s}o_{i,\ell}\nabla\cdot\bm{\rho}^{(i)},\sum_{j=1}^{s}o_{j,\ell}\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{\ell=1}^{s}\lambda_{\ell}\left\|\sum_{i=1}^{s}o_{i,\ell}\nabla\cdot\bm{\rho}^{(i)}\right\|^{2}\geq 0.\end{array}$
@@
(56b) and the normal component of (56c) imply $\mathbf{n}\cdot\bm{\rho}^{(i)}=-\lambda\mathbf{n}\cdot\mathbf{w}^{(i)}$. Then arguments similar to those for (58) yield (59). ∎
Corollary 2.
@@
Suppose the RK method employed in the GePUP-ES-RK scheme (56)–(55) is algebraically stable in the sense of Definition 2. Then (40) implies
@@
Proof.
The initial condition (40) states $\nabla\cdot\mathbf{w}(t_{0})=\nabla\cdot\mathbf{u}(t_{0})=0$ and $\mathbf{n}\cdot\mathbf{w}(t_{0})|_{\partial\Omega}=\mathbf{n}\cdot\mathbf{u}(t_{0})|_{\partial\Omega}=0$. Then the first clauses of (62) and (65) follow from Theorem 9 and (ABS-1) in Definition 2. (58) dictates $\|\nabla(\nabla\cdot\mathbf{w}^{(i)})\|=0$ for each stage. Then the second clause of (62) follows from the boundary condition $\nabla\cdot\mathbf{w}|_{\partial\Omega}=0$ in (39b). Similarly, (59) dictates the second clause of (65). ∎
Theorem 10.
Suppose that the body force $\mathbf{g}$ in the GePUP-E formulation is conservative, that the initial condition of $\mathbf{w}$ is (40), and that the employed RK method in (56)–(55) is algebraically stable in the sense of Definition 2. Then the energy dissipation of the GePUP-ES-RK scheme (56)–(55) is governed by
@@
Proof.
Denote $\bm{\sigma}^{(i)}:=\mathscr{P}\bm{\rho}^{(i)}$ and we have from (56c,d) and (55)
$\begin{array}[]{c}\forall i=1,\ldots,s,\ \ \mathbf{u}^{(i)}=\mathscr{P}\mathbf{w}^{(i)}=\mathscr{P}\left(\mathbf{w}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\rho}^{(j)}\right)=\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\sigma}^{(j)};\\ \mathbf{u}^{n+1}=\mathscr{P}\mathbf{w}^{n+1}=\mathscr{P}\left(\mathbf{w}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\rho}^{(i)}\right)=\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\sigma}^{(i)}.\end{array}$
@@
where the last step follows from (38). Then (ABS-2) in Definition 2 yields
@@
Write $\alpha_{i}:=I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)})=\left\langle\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}\right\rangle$ for each stage $i=1,\ldots,s$. Then
@@
where the first equality follows from $\bm{\sigma}^{(i)}=\mathscr{P}\bm{\rho}^{(i)}$, the second from (26), (56g,h), and Lemma 9, the third from (56c) and the definition of $\alpha_{i}$, and the last from $\mathbf{g}$ being conservative, Lemma 9, (53), and Corollary 2.
Since the Leray-Helmholtz projection has no control over the tangential velocity, in (56g) we would have $\bm{\tau}\cdot\mathbf{u}^{(i)}\neq 0$ if $\nabla\cdot\mathbf{w}^{(i)}\neq 0$. Fortunately Corollary 2 dictates $\nabla\cdot\mathbf{w}^{(i)}=0$ and $\mathbf{n}\cdot\mathbf{w}^{(i)}|_{\partial\Omega}=0$, then the Leray-Helmholtz projection in (56g) reduces to the identity. Thus the second integral in the last line of (53) vanishes in GePUP-ES-RK. A related observation is that, although $\nabla q$ is orthogonal to $\mathbf{w}$, it cannot be arbitrary as it must make $\bm{\rho}^{(j)}$ in (56d) divergence-free; otherwise the Leray-Helmholtz projection in (56g) would not reduce to the identity.
Substitute $\left\langle\mathbf{u}^{(i)},\bm{\sigma}^{(i)}\right\rangle=-r^{(i)}\alpha_{i}-\nu\left\|\nabla\mathbf{u}^{(i)}\right\|^{2}$ into ()(*)( ∗ ) and we have
@@
Similarly, the positive semi-definiteness of $M$ in Definition 2 yields
@@
Except on staggered grids, the initial condition of $\mathbf{w}$ being incompressible in (40) cannot be exactly fulfilled in practical computations. However, the mechanism of divergence decay ensures that $\mathbf{w}$ converge to $\mathbf{u}$ sufficiently fast; see Theorem 2. This is also related to the millennium problem on the well-posedness of the INSE. If the solution of the INSE blows up, then its divergence must blow up first. Therefore, our computation only works when the INSE admits a bounded solution. After all, one cannot expect to solve the millennium problem by reformulating INSE.

### v2

@@ A GePUP-ES-RK scheme for solving INSE with no-slip conditions is a semidiscrete algorithm of the form
$\mathbf{w}^{(i)}$ $=\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\rho}^{(j)}$ $\text{in }\Omega,$ (49d)
$\mathbf{w}^{(i)}\cdot\bm{\tau}$ $=0,~\nabla\cdot\mathbf{w}^{(i)}=0$ $\text{on }\partial\Omega,$ (49e)
@@
$\mathbf{u}^{(i)}$ $=\mathscr{P}\mathbf{w}^{(i)}$ $\text{in }\Omega,$ (49g)
$\mathbf{n}\cdot\mathbf{u}^{(i)}$ $=0$ $\text{on }\partial\Omega.$ (49h)
As suggested by (49d) and the first equation of (48), $\mathbf{u}^{n}$ is used as the initial condition of $\mathbf{w}$ for time integration within the interval $t^{n},t^{n+1}]$. However, it is emphasized that we do not write $\mathbf{w}^{n}=\mathbf{u}^{n}$ because this would cause a notation clash to $\mathbf{w}^{n+1}$, which, by (49), is the evolutionary velocity before the final projection and thus needs not be divergence free. In Definition 5.1, the approximations $\widetilde{\mathbf{u}}^{(i)}$ in $\Omega$ may be obtained by polynomial interpolation based on stage values of recent time steps. In this work, we fit a cubic polynomial $\mathbf{p}$ from the known velocity $\hat{\mathbf{u}}_{n}^{(j)}$ at time instances $t^{n}+\hat{c}_{j}k$ with $j=0,1,2,3$ and then approximate $\mathbf{u}(t^{n}+c_{i}k)$ with
@@
  1. (EAV-1)
    for the first two time steps $n=0,1$, we calculate $\hat{\mathbf{u}}^{(j)}$ by GePUP-ERK zhang16:_GePUP , the explicit RK method for solving the GePUP formulation, with time step size $\frac{1}{3}k$ and then fit $\mathbf{p}$ in (50) with $\left[\hat{c}_{0},\hat{c}_{1},\hat{c}_{2},\hat{c}_{3}\right]=\left[0,\frac{1}{3},\frac{2}{3},1\right]$;
  2. (EAV-2)
@@
$t^{n}$$t^{n+1}$$\widetilde{\mathbf{u}}_{n}^{\left(1\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(2\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(3\right)}$$t^{n}+c_{1}k$$t^{n}+c_{2}k$$t^{n}+c_{3}k$$t^{n}$$t^{n}+\frac{1}{3}k$$t^{n}+\frac{3}{3}k=t^{n+1}$$\mathbf{u}^{n}$$\mathbf{u}^{n+1}$$\hat{\mathbf{u}}_{n}^{\left(0\right)}=\mathbf{u}^{n}$$\hat{\mathbf{u}}_{n}^{\left(1\right)}$$\hat{\mathbf{u}}_{n}^{\left(3\right)}$$t^{n}+\frac{2}{3}k$$\hat{\mathbf{u}}_{n}^{\left(2\right)}$
(a) (EAV-1) for $n=0,1$
$t^{n}$$t^{n+1}$$t^{n-2}$$\widetilde{\mathbf{u}}_{n}^{\left(1\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(2\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(3\right)}$$t^{n}+c_{1}k$$t^{n}+c_{2}k$$t^{n}+c_{3}k$$t^{n}$$t^{n-1}$$t^{n+1}=t^{n}+k$$\mathbf{u}^{n}$$\mathbf{u}^{n+1}$$\hat{\mathbf{u}}_{n}^{\left(2\right)}=\mathbf{u}^{n}$$\hat{\mathbf{u}}_{n}^{\left(1\right)}=\mathbf{u}^{n-1}$$\hat{\mathbf{u}}_{n}^{\left(3\right)}$$\hat{\mathbf{u}}_{n}^{\left(0\right)}=\mathbf{u}^{n-2}$
(b) (EAV-2) for $n\geq 2$
Figure 1: Estimating the auxiliary velocity $\widetilde{\mathbf{u}}^{(i)}$ (EAV) in GePUP-ES-RK by (50).
Theorem 5.2
Suppose the RK method employed in the GePUP-ES-RK scheme (48) is algebraically stable in the sense of Definition 2.9. Then (34) implies
@@
$\forall n\!\in\!\mathbb{N},\ \forall\mathbf{x}\!\in\!\partial\Omega,~\left|\mathbf{n}\!\cdot\!\mathbf{w}^{n+1}(\mathbf{x})\right|^{2}\!-\!\left|\mathbf{n}\!\cdot\!\mathbf{w}^{n}(\mathbf{x})\right|^{2}\!\leq\!-2k\lambda\sum_{i=1}^{s}b_{i}\left|\mathbf{n}\!\cdot\!\mathbf{w}^{(i)}(\mathbf{x})\right|^{2}\!.$ (52)
Proof
Take divergence of (49c), apply (49a) and the commutativity of $\Delta$ and $\nabla\cdot$, and we have $\nabla\cdot\bm{\rho}^{(i)}=\nu\Delta\left(\nabla\cdot\mathbf{w}^{(i)}\right)$. Then
@@
$\begin{array}[]{rl}&\left\langle\nabla\cdot\mathbf{w}^{n+1},\nabla\cdot\mathbf{w}^{n+1}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\mathbf{u}^{n}+k\sum_{j=1}^{s}b_{j}\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\mathbf{u}^{n}\right\rangle+k\sum_{i=1}^{s}b_{i}\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\bm{\rho}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\bm{\rho}^{(j)}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\mathbf{u}^{n}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ &+k\sum_{i=1}^{s}b_{i}\left\langle\nabla\cdot\mathbf{w}^{(i)}-k\sum_{j=1}^{s}a_{i,j}\nabla\cdot\bm{\rho}^{(j)},\nabla\cdot\bm{\rho}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\nabla\cdot\mathbf{w}^{(j)}-k\sum_{i=1}^{s}a_{j,i}\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\mathbf{u}^{n}\right\rangle-2\nu k\sum_{i=1}^{s}b_{i}\left\|\nabla\left(\nabla\cdot\mathbf{w}^{(i)}\right)\right\|^{2}\\ &-k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle,\end{array}$
@@
where the last step follows from (31), (28), and (49e). By Definition 2.9, the algebraic stability implies the symmetric positive semi-definiteness of $M$ and thus we have $M=O\Lambda O^{T}$, where $O$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix with $\lambda_{\ell}=\Lambda_{\ell,\ell}\geq 0$ for each $\ell=1,\ldots,s$. Hence (51) follows from
@@
$\begin{array}[]{rl}&\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{i=1}^{s}\sum_{j=1}^{s}\left(\sum_{\ell=1}^{s}o_{i,\ell}\lambda_{\ell}o_{j,\ell}\right)\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{\ell=1}^{s}\lambda_{\ell}\left\langle\sum_{i=1}^{s}o_{i,\ell}\nabla\cdot\bm{\rho}^{(i)},\sum_{j=1}^{s}o_{j,\ell}\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{\ell=1}^{s}\lambda_{\ell}\left\|\sum_{i=1}^{s}o_{i,\ell}\nabla\cdot\bm{\rho}^{(i)}\right\|^{2}\geq 0\end{array}$
and the fact that the Leray-Helmholtz projection $\mathscr{P}$} do not increase the divergence of velocity, i.e., $\|\nabla\cdot\mathbf{u}^{n}\|^{2}=\|\nabla\cdot{\mathscr{P}}\mathbf{w}^{n}\|^{2}\leq\|\nabla\cdot\mathbf{w}^{n}\|^{2}$.
@@
(49b) and the normal component of (49c) imply $\mathbf{n}\cdot\bm{\rho}^{(i)}=-\lambda\mathbf{n}\cdot\mathbf{w}^{(i)}$. Then (52) follows from arguments similar to those for (51) and
$\forall n\in\mathbb{N},\ \forall\mathbf{x}\in\partial\Omega,\ \left|\mathbf{n}\cdot\mathbf{u}^{n}(\mathbf{x})\right|^{2}-\left|\mathbf{n}\cdot\mathbf{w}^{n}(\mathbf{x})\right|^{2}=-\left|\mathbf{n}\cdot\mathbf{w}^{n}(\mathbf{x})\right|^{2}\leq 0.\hskip 22.76219pt\qed$
In the fully discrete sense, we discretize each continuous operator in (51) add the corresponding discretization error, and obtain a discrete version of (51). As discussed in Section 6, this discrete version of (51) is helpful in understanding the evolution of the discrete velocity divergence.
Interestingly, (51) is useless for first-order finite difference/volume methods. Suppose $\mathbf{w}^{n}=\mathbf{w}(t^{n})+O(h^{2})$. Then, due to the factor $k$ and the second derivative on the RHS of (51), the discrete version of (51) has an $O(1)$ error. In contrast, a fourth-order method has $\mathbf{w}^{n}=\mathbf{w}(t^{n})+O(h^{4})$. With the added discretization error approaching zero as $h,k\rightarrow 0$, the discrete version of (51) indeed has control over the evolution of the discrete velocity divergence. The above discussion suggests a crucial advantage of fourth-order methods over first-order methods.
Corollary 5.3
@@
Suppose the RK method employed in the GePUP-ES-RK scheme (48) is algebraically stable in the sense of Definition 2.9. Then the initial condition (33) implies
@@
Proof
(33) gives $\nabla\cdot\mathbf{w}(t_{0})=\nabla\cdot\mathbf{u}(t_{0})=0$ and $\mathbf{n}\cdot\mathbf{w}(t_{0})|_{\partial\Omega}=\mathbf{n}\cdot\mathbf{u}(t_{0})|_{\partial\Omega}=0$. Then the first clauses of (55) and (58) follow from Theorem 5.2 and the first condition in Definition 2.9. (51) dictates $\|\nabla(\nabla\cdot\mathbf{w}^{(i)})\|=0$ for each stage. Then the second clause of (55) follows from the boundary condition $\nabla\cdot\mathbf{w}|_{\partial\Omega}=0$ in (32b). Similarly, (52) dictates the second clause of (58). ∎
Of course Corollary 5.3 holds only in the semi-discrete sense. Since Corollary 5.3 is used in proving Theorem 5.4, the decay of the modified energy in the fully discrete sense depends on that of the discrete velocity divergence.
Theorem 5.4
Suppose that the body force $\mathbf{g}$ in the GePUP-E formulation is conservative, that the initial condition of $\mathbf{w}$ is (33), and that the employed RK method in (48) is algebraically stable in the sense of Definition 2.9. Then the energy dissipation of the GePUP-ES-RK scheme (48) is governed by
@@
Proof
Denote $\bm{\sigma}^{(i)}:=\mathscr{P}\bm{\rho}^{(i)}$ and we have from (49c,d) and (48)
$\begin{array}[]{c}\forall i=1,\ldots,s,\ \ \mathbf{u}^{(i)}=\mathscr{P}\mathbf{w}^{(i)}=\mathscr{P}\left(\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\rho}^{(j)}\right)=\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\sigma}^{(j)};\\ \mathbf{u}^{n+1}=\mathscr{P}\mathbf{w}^{n+1}=\mathscr{P}\left(\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\rho}^{(i)}\right)=\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\sigma}^{(i)}.\end{array}$
@@
where the last step is due to (31). The second condition in Definition 2.9 gives
@@
Write $\alpha_{i}:=I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)})=\left\langle\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}\right\rangle$ for each $i=1,\ldots,s$. Then
@@
where the first equality follows from $\bm{\sigma}^{(i)}=\mathscr{P}\bm{\rho}^{(i)}$, the second from (19), (49g,h), and Lemma 3.10, the third from (49c) and the definition of $\alpha_{i}$, and the last from $\mathbf{g}$ being conservative, Lemma 3.10, (46), and Corollary 5.3.
Since the Leray-Helmholtz projection $\mathscr{P}$ has no control over the tangential velocity, in (49g) we would have $\bm{\tau}\cdot\mathbf{u}^{(i)}\neq 0$ if $\nabla\cdot\mathbf{w}^{(i)}\neq 0$. Fortunately Corollary 5.3 dictates $\nabla\cdot\mathbf{w}^{(i)}=0$ and $\mathbf{n}\cdot\mathbf{w}^{(i)}|_{\partial\Omega}=0$, then in (49g) $\mathscr{P}$ reduces to the identity. Thus the second integral in the last line of (46) vanishes in GePUP-ES-RK. A related observation is that, although $\nabla q$ is orthogonal to $\mathbf{w}$, it cannot be arbitrary as it must make $\bm{\rho}^{(j)}$ in (49d) divergence-free; otherwise in (49g) $\mathscr{P}$ would not reduce to the identity.
Substitute $\left\langle\mathbf{u}^{(i)},\bm{\sigma}^{(i)}\right\rangle=-r^{(i)}\alpha_{i}-\nu\left\|\nabla\mathbf{u}^{(i)}\right\|^{2}$ into ()(*) and we have
@@
Similarly, the positive semi-definiteness of $M$ in Definition 2.9 yields
@@
Except on staggered grids, the initial condition of $\mathbf{w}$ being incompressible in (33) cannot be exactly fulfilled in practical computations. However, the mechanism of divergence decay ensures that $\mathbf{w}$ converge to $\mathbf{u}$ sufficiently fast; see Theorem 5.3. This is also related to the millennium problem on the well-posedness of the INSE. If the solution of the INSE blows up, then its divergence must blow up first. Therefore, our computation only works when the INSE admits a bounded solution. After all, one cannot expect to solve the millennium problem by reformulating INSE.